OpenStudy (anonymous):

Determine the symmetry of the function y=-x^3 (even, odd, or no symmetry) and explain how you arrived at your answer. Again, this wasn't in my lesson... Unless I missed it, trying to find it again...

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

So it would be no sym. Right?

OpenStudy (unklerhaukus):

looks like it has rotational symmetry to me

OpenStudy (unklerhaukus):

a function \(f(x)\) has odd symmetry iff \[f(-x)=-f(x)\] even symmetry iff \[f(-x)=f(x)\]

OpenStudy (unklerhaukus):

\[y=f(x)=-x^3\] \[f(-x)=-(-x)^3=\]